# Unleashing the Potential of Tracklets for Unsupervised Video Person Re-Identification

### v1

@@ IV-D Ablation Studies
Influence of different pooling strategies. GeM pooling generalizes max and average pooling and is more robust to noises in the unsupervised scenario, so it is adopted in SSR-C to obtain the feature embedding of each frame. We compare different pooling strategies in Tab. VI and find GeM more effective. Note that by adopting either GAP or GMP, our method still surpasses state-of-the-art methods on both datasets by a large margin, showing its effectiveness and robustness.
@@
Analysis of $l$. Tracklet partition stride $l$ controls the granularity of sub-tracklet representation. As shown in Fig. 4 (a), large $l$ exacerbates variations within tracklets and contributes to inferior performance, which confirms our motivation. Conversely, small $l$ brings redundancy and a massive data column. For a tradeoff between performance and efficiency, we set $l=32$.
@@
Analysis of $K$. The value of $K$ in our CSC loss is dynamically determined considering different numbers of positive sub-tracklets for different people. Besides dynamicity, CSC loss guides learning at the fine-grained sub-tracklet level. We compare with methods using a fixed $K$ design [25, 37] in Tab. I and SSR-C shows great superiority. We also try manually setting $K$ with different values in Fig. 4 (b). They all perform worse and the optimal $K$ needs cherry-pick.
@@
Analysis of $\lambda$. As stated in Eq. 10, $\lambda$ controls the weight of smoothing assigned to $K$ positive sub-clusters. When $\lambda=0$, the model does not consider other positive sub-tracklets from the same tracklet or different tracklets with the same identity. Since it pushes different sub-clusters with the same identity away, it achieves sub-optimal results as shown in Fig. 4 (c). Compared to the widely used InfoNCE loss, CSC loss shows better performance for different choices of $\lambda$. As the value of $\lambda$ increases, higher weights are assigned to other potential positive sub-clusters, which would distract the model learning, so we observe a downward trend in performance in Fig. 4 (c). The best results are achieved when $\lambda=0.1$.
@@
Figure 5: (a) Influence of different values of $\delta$ on the performance, and (b) the average number of filtered frames at different epochs with various values of $\delta$. The ablation is performed on Duke-V and the optimal value are directly applied to MARS without further tuning.
@@
Analysis of $\delta$. We try various values of $\delta$ for noise filtering in Fig. 5 (a), and note an overall performance improvement. According to Eq. 7, as the value of $\delta$ increases, more frames would be identified as noises and filtered. When $\delta$ is relatively small, some noisy frames cannot be filtered, which would misguide clustering and result in sub-optimal results. As $\delta$ is overly large, the model also achieves inferior results, since more frames are identified as noises and some frames that contain useful information may be filtered excessively. In our experiments, noises within tracklets can be effectively filtered with $\delta=0.7$.
@@
Amount of filtered noisy frames. The results in Fig. 5 (b) support our deduction that more frames can be filtered as $\delta$ increases. According to Fig. 5 (a), the optimal performance is achieved when setting $\delta$ to 0.7, which filters about an average of 30 frames per tracklet. As shown in Fig. 5 (b), we also observe that even when the value of $\delta$ remains constant, the number of filtered noisy frames varies dynamically as the training progresses. When $\delta$ is less than 0.6, more noisy frames are filtered in the later training stage. However, when the value of $\delta$ is larger than 0.6, the number of filtered noisy frames gradually decreases along with the training process. Intuitively, relatively easier frames should be used for model learning at first, and as the feature extraction capability of the model improves, some moderately hard frames can be incorporated to enhance the robustness of the model. When $\delta=0.7$, which is the value used in our experiments, the trend of the number of filtered noisy frames matches the above expectation.
@@
Figure 6: The statistical information of clustering. The number of correct clusters, cross-camera clusters, and incorrect clusters using “Baseline $+$ NFTP” and our proposed method are displayed, respectively. Cross-camera clusters are a subset of correct clusters. For each dataset, the ground truth number of identities is shown in the upper right.
@@
Statistics of clustering. To intuitively show why our proposed SSR-C can better guide model learning, we also compare its statistics of clustering with “Baseline $+$ NFTP” (Since without NFTP, the model would not converge on Duke-V as discussed in Sec. IV-D). As shown in Fig. 6, our proposed SSR-C contributes to more correct clusters on both datasets. What’s more, it exhibits significantly fewer incorrect clusters. We also compare the number of clusters with cross camera samples, i.e., cross-camera clusters. The more correct cross-camera clusters of SSR-C indicate that our model can learn richer information from multiple camera views and further enable cross-camera identity matching.
V Visualizations
@@
Figure 9: Visualization of top eight retrieval results of our proposed SSR-C and “Baseline + NFTP” on MARS, respectively. The correct and wrong results are in green and red boxes, respectively. To simplify the presentation, we showcase only the first frame of each tracklet.
Visualization of noisy frames. The presence of noisy frames can render each tracklet feature non-discriminative, misleading the clustering algorithm to generate incorrect pseudo labels and disturbing feature learning. Especially when we apply our proposed tracklet partition, sub-tracklets with noises may cause relay errors in the following sub-cluster merging step. Therefore, noise filtering plays a pivotal role in SSR-C. To intuitively understand how it works and show its effectiveness, we visualize the filtered noisy frames at different training stages in Fig. 7. As shown in Fig. 7 (a) and (b), there may be occlusions caused by other pedestrians or objects. If we perform clustering based on the unfiltered tracklet features, tracklets belonging to different individuals may be close in the embedding space and be wrongly clustered. Similarly, frames with great scale changes and background interference can also be regarded as noises and be filtered, as shown in Fig. 7 (c).
What’s more, some slightly noisy frames are discarded during the early training stages, but are included for training later. For example, as shown in the first row of Fig. 7 (a) and (b), there are heavy occlusions from other pedestrians or objects, thereby only limited identity information remains in them. Since there is no supervision of ground-truth identities, these noisy frames can interfere with the tracklet features, which in turn interferes with reliable clustering and disrupts the model optimization. Nevertheless, as shown in the second row of Fig. 7 (a) and (b), some noisy frames can still contain useful identity information. With moderate noises, they can help the model deal with occlusion interference in the real world. Therefore, thanks to our dynamic noise-filtering design, they can be used for robust feature learning. Another example is shown in Fig. 7 (c), where with the pedestrian walking into the room and away from the camera, the scale varies and the background interference increases. As the training progresses, some difficult noisy frames (in the first row) are filtered all the way, while some moderately challenging frames (in the second row) are gradually used for subsequent model optimization. Since the model’s capability and robustness improve as training, gradually introducing relatively hard frames can improve its robustness in dealing with various interference.
@@
Clustering results. We also use t-SNE [61] to visualize the clustering results in Fig. 8. For the baseline based on tracklets, since there are noises within tracklets, different identities are aggregated together incorrectly at the early stage in Fig. 8 (a). When it comes to the later training stage, the clustering results are still unsatisfactory with a lot of wrong clusters. In contrast, for SSR-C, multiple sub-tracklets from the same tracklet, acting as intra-class cross-tracklet relays, can effectively aggregate various cross-camera tracklets of the same identity as shown at the bottom of Fig. 8 (b). Additionally, although some tracklets with different identities are wrongly clustered in the early training stage, they are separated in the later training stage. By comparing the clustering results at the late training stage in Fig. 8, we can observe that SSR-C contributes to better intra-class compactness and inter-class separability. Therefore, SSR-C can generate more accurate pseudo labels and promote the feature extraction capability of the model.
@@
Retrieval results. To intuitively demonstrate the effectiveness of our method, we visualize the retrieval results in Fig. 9. Considering the poor performance of the baseline, we compare the retrieval results between our proposed SSR-C and “Baseline $+$ NFTP”. As shown in Fig. 9, SSR-C can better encourage the model to learn useful identity information. For example, “Baseline $+$ NFTP” may be influenced by similar color appearance information as shown in Fig. 9 (a), while our method can deal with it well to some extent. The results of (b) and (d) in Fig. 9 show that our proposed SSR-C is more capable of attending to the effective regions of pedestrians, and (c) and (d) in Fig. 9 demonstrate its effectiveness in handling view changes and scene variations. Due to poor clustering accuracy, “Baseline $+$ NFTP” cannot handle complicated variations and visual interference, leading to wrong matches. Overall, the results validate the heightened robustness of our proposed SSR-C across diverse challenging scenarios, including lighting variations, scene changes, viewpoint changes, similar appearances and misalignment induced by tracking. Moreover, even without relying on extra information, SSR-C demonstrates improved feature learning and retrieval results.
VI Conclusion
@@
To overcome the dependence on identity labels and exploit rich information naturally in tracklets, we propose the Self-Supervised Refined Clustering (SSR-C) framework for unsupervised video person Re-ID. It unleashes the potential of tracklets for effective model optimization. Specifically, we first design the NFTP module to eliminate the noises in tracklets and partition each tracklet into multiple sub-tracklets. With noise-filtered sub-tracklets, we take advantage of the self-supervised signal that the sub-tracklets from the same tracklet indicate a consistent identity, and propose a progressive merging strategy to achieve reliable clustering and generate accurate pseudo labels. To exploit the associations between sub-tracklets and facilitate feature learning, we further propose CSC loss. Without relying on manual annotations or auxiliary information, SSR-C approaches advanced supervised methods in performance. There is no additional overhead during inference, making it efficient in real-world applications. Extensive experiments demonstrate its effectiveness and superiority. In the future, it is promising to investigate different partition strategies to address variations within long tracklets more effectively and filter noisy frames dynamically according to the model’s ability at different training stages.

### v2

@@ IV-E6 Influence of Different Pooling Strategies
Figure 4: Influence of different values of (a) $l$, (b) $K$, (c) $\lambda$. The blue and red dashed lines in (b) denote the Rank-1 and mAP accuracies of our proposed method, respectively. The ablation experiments are performed on the Duke-V dataset and the optimal hyper-parameter values are directly applied to MARS.
IV-E7 Analysis of $l$
@@
Tracklet partitioning stride $l$ controls the granularity of sub-tracklet representation. As shown in Fig. 4 (a), large $l$ exacerbates variations within tracklets and contributes to inferior performance, which confirms our motivation. Conversely, small $l$ brings redundancy and a massive data column. For a tradeoff between performance and efficiency, we set $l=32$.
IV-E8 Analysis of $K$
@@
The value of $K$ in our CSC loss is dynamically determined considering different numbers of positive sub-tracklets for different people. Besides dynamicity, CSC loss guides learning at the fine-grained sub-tracklet level. We compare with methods using a fixed $K$ design [27, 39] in Tab. I and SSR-C shows great superiority. We also try manually setting $K$ with different values in Fig. 4 (b). They all perform worse and the optimal $K$ needs cherry-pick.
IV-E9 Analysis of $\lambda$
@@
As stated in Eq. 11, $\lambda$ controls the weight of smoothing assigned to $K$ positive sub-clusters. When $\lambda=0$, the model does not consider other positive sub-tracklets from the same tracklet or different tracklets with the same identity. Since it pushes different sub-clusters with the same identity away, it achieves sub-optimal results as shown in Fig. 4 (c). Compared to the widely used InfoNCE loss, CSC loss shows better performance for different choices of $\lambda$. As the value of $\lambda$ increases, higher weights are assigned to other potential positive sub-clusters, which would distract the model learning, so we observe a downward trend in performance in Fig. 4 (c). The best results are achieved when $\lambda=0.1$.
@@
Figure 5: (a) Influence of different values of $\delta$ on the performance, and (b) the average number of filtered frames at different epochs with various values of $\delta$. The blue and red dashed lines in (a) denote the Rank-1 and mAP accuracies of the variant without using NF, respectively. The ablation is performed on Duke-V and the optimal value are directly applied to MARS without further tuning.
IV-E10 Analysis of $\delta$
@@
We try various values of $\delta$ for noise filtering in Fig. 5 (a), and note an overall performance improvement when applying NF. According to Eq. 8, as the value of $\delta$ increases, more frames would be identified as noises and filtered. When $\delta$ is relatively small, some noisy frames cannot be filtered, which would misguide clustering and result in sub-optimal results. As $\delta$ is overly large, the model also achieves inferior results, since more frames are identified as noises and some frames that contain useful information may be filtered excessively. In our experiments, noises within tracklets can be effectively filtered with $\delta=0.7$.
IV-E11 Amount of Filtered Noisy Frames
@@
The results in Fig. 5 (b) support our deduction that more frames can be filtered as $\delta$ increases. We also observe that even when $\delta$ remains constant, the number of filtered noisy frames varies dynamically as the training progresses. When $\delta$ is less than 0.5, more noisy frames are filtered in the later training stage. However, when the value of $\delta$ is larger than 0.5, the number of filtered noisy frames gradually decreases along with the training process. Intuitively, relatively easier frames should be used for model learning at first, and as the feature extraction capability of the model improves, some moderately hard frames can be incorporated to enhance the robustness of the model. When $\delta=0.7$, which is the value we use, the trend of the number of filtered noisy frames matches the above expectation.
@@
Figure 6: The statistical information of clustering. The number of correct clusters, cross-camera clusters, and incorrect clusters using “Baseline ++ NFTP” and our method are displayed, respectively. For correct clusters, all the samples within the cluster share the same identity label. Cross-camera clusters are a subset of correct clusters. For each dataset, the ground truth number of identities is shown in the upper right.
IV-E12 Statistics of Clustering
@@
To intuitively show why SSR-C can better guide model learning, we also compare its statistics of clustering with “Baseline ++ NFTP” (Since without NFTP, the model would not converge on Duke-V as discussed in Sec. IV-E). As shown in Fig. 6, SSR-C contributes to a larger proportion of correct clusters on both datasets. It also exhibits significantly fewer incorrect clusters, which demonstrates the efficacy of SSR-C in improving clustering results. We also compare the number of clusters with cross-camera samples, i.e., cross-camera clusters. The more correct cross-camera clusters of SSR-C can help our model learn richer information from multiple camera views and promote cross-camera identity matching.
TABLE VII: Ablation studies of different clustering algorithms. The experiments are conducted on the Duke-V dataset. “#Clusters” denotes the preset number of clusters.
Clustering Alg. #Clusters mAP Rank-1
K-means 500 7.8 11.8
600 21.5 26.2
700 93.0 93.7
800 93.0 94.2
900 93.3 93.3
1000 94.7 94.9
1100 93.3 93.9
1200 91.5 92.9
DBSCAN Dynamic 95.3 95.9
IV-E13 Ablation of the Clustering Algorithm
To show the effectiveness of applying DBSCAN [51] for clustering in our proposed SSR-C framework, we also try replacing it with K-means [43], which needs presetting the number of clusters. As shown in Tab. VII, when the number of clusters is too small, different identities are incorrectly grouped together, weakening the model’s ability to distinguish between individuals. While increasing the number of clusters can mitigate this to some extent, the fixed nature of K-means clustering causes it to push away features from the same identity in the whole training process, harming discriminative learning. Additionally, searching for the optimal number of clusters requires prior information of the data and many preliminary experiments, making K-means inflexible and unrealistic in applications. As shown in Tab. VII, when the number of clusters is not appropriately set (e.g., 500 and 600), the performance degrades significantly. In contrast, DBSCAN dynamically adjusts the number of clusters based on the underlying feature distribution, allowing for more accurate and adaptive clustering. Combined with our proposed self-supervised refined clustering design, this flexibility enables the model to continuously refine clustering results in response to improved feature representations, leading to better performance in the unsupervised setting. The results show that our SSR-C based on DBSCAN is capable of generating reliable pseudo labels to achieve effective feature learning.
IV-F Visualizations
@@
IV-F1 Visualization of Noisy Frames
The presence of noisy frames can render each tracklet feature non-discriminative, misleading the clustering algorithm to generate incorrect pseudo labels and disturbing feature learning. Especially when we apply our proposed tracklet partition, sub-tracklets with noises may cause relay errors in the following sub-cluster merging step. Therefore, noise filtering plays a pivotal role in SSR-C. To intuitively understand how it works and show its effectiveness, we visualize the filtered noisy frames at different training stages in Fig. 7. As shown in Fig. 7 (a) and (b), there may be occlusions caused by other pedestrians or objects. If we perform clustering based on the unfiltered tracklet features, tracklets belonging to different individuals may be close in the embedding space and be wrongly clustered. Similarly, frames with great scale changes and background interference can also be regarded as noises and be filtered, as shown in Fig. 7 (c).
Additionally, some slightly noisy frames are discarded during the early training stages, but are included for training later. For example, as shown in the first row of Fig. 7 (a) and (b), there are heavy occlusions from other pedestrians or objects, thereby only limited identity information remains in them. Since there is no supervision of ground-truth identities, these noisy frames can interfere with the tracklet features, which in turn interferes with reliable clustering and disrupts the model optimization. Nevertheless, as shown in the second row of Fig. 7 (a) and (b), some noisy frames can still contain useful identity information. With moderate noises, they can help the model deal with occlusion interference in the real world. Therefore, thanks to our dynamic noise-filtering design, they can be used for robust feature learning. Another example is shown in Fig. 7 (c), where with the pedestrian walking into the room and away from the camera, the scale varies and the background interference increases. As the training progresses, some difficult noisy frames (in the first row) are filtered all the way, while some moderately challenging frames (in the second row) are gradually used for subsequent model optimization. Since the model’s capability and robustness improve as training, gradually introducing relatively hard frames can improve its robustness in dealing with various interference.
@@ IV-F1 Visualization of Noisy Frames
IV-F2 Clustering Results
@@
We also use t-SNE [74] to visualize the clustering results in Fig. 8. For the baseline based on tracklets, since there are noises within tracklets, different identities are aggregated together incorrectly at the early stage in Fig. 8 (a). When it comes to the later training stage, the clustering results are still unsatisfactory with a lot of wrong clusters. In contrast, for SSR-C, multiple sub-tracklets from the same tracklet, acting as intra-class cross-tracklet relays, can effectively aggregate various cross-camera tracklets of the same identity as shown at the bottom of Fig. 8 (b). Additionally, although some tracklets with different identities are wrongly clustered in the early training stage, they are separated in the later training stage. By comparing the clustering results at the late training stage in Fig. 8, we can observe that SSR-C contributes to better intra-class compactness and inter-class separability. Therefore, SSR-C can generate more accurate pseudo labels and promote the feature extraction capability of the model.
Figure 9: Visualization of top eight retrieval results of our proposed SSR-C and “Baseline + NFTP” on MARS, respectively. The correct and wrong results are in green and red boxes, respectively. To simplify the presentation, we showcase only the first frame of each tracklet. Best viewed in color and zoomed in.
IV-F3 Retrieval Results
@@
To intuitively demonstrate the effectiveness of our method, we visualize the retrieval results in Fig. 9. Considering the poor performance of the baseline, we compare the retrieval results between our proposed SSR-C and “Baseline ++ NFTP”. As shown in Fig. 9 (a), SSR-C can better encourage the model to learn useful identity information, while “Baseline ++ NFTP” may be influenced by similar color appearance information. The results of (b) and (c) show that our SSR-C is more capable of attending to the effective regions of pedestrians, and the results of (c) and (d) demonstrate its effectiveness in handling view changes and scene variations. Due to poor clustering accuracy, “Baseline ++ NFTP” cannot handle complicated variations and visual interference, leading to wrong matches. Overall, the results validate the heightened robustness of our proposed SSR-C across diverse challenging scenarios, including lighting variations, scene changes, viewpoint changes, similar appearances, and misalignment induced by tracking. We also observe some failure cases of SSR-C in (e) and (f), where people with similar body shapes wear similar clothes or carry similar backpacks. It is expected to improve its robustness by incorporating other auxiliary information such as attributes and part-based modules. Without relying on extra information, SSR-C demonstrates improved feature learning and retrieval results.
V Conclusion and Outlook
@@
To overcome the dependence on identity labels and exploit rich information naturally in tracklets, we propose the Self-Supervised Refined Clustering (SSR-C) framework for unsupervised video person Re-ID. It unleashes the potential of tracklets for effective model optimization. Specifically, we first design the NFTP module to eliminate the noises in tracklets and partition each tracklet into multiple sub-tracklets. With noise-filtered sub-tracklets, we take advantage of the self-supervised signal that the sub-tracklets from the same tracklet indicate a consistent identity, and propose a progressive sub-cluster merging strategy to achieve reliable clustering and generate accurate pseudo labels. To exploit the associations between sub-tracklets and facilitate feature learning, we further propose CSC loss. Extensive experiments demonstrate the effectiveness and superiority of our proposed SSR-C. Without relying on manual annotations or auxiliary information, SSR-C approaches advanced supervised methods in performance. There is no additional overhead during inference, making it efficient in real-world applications.
In our SSR-C framework, we propose a straightforward strategy for tracklet partition, i.e., sequentially dividing each noise-filtered tracklet into multiple sub-tracklets. It would be promising to explore alternative partitioning strategies in future work, to better handle the variations present in long tracklets. Moreover, as illustrated in Figs. 5 (b) and 7, the number of filtered noisy frames varies throughout the training process. Although our noise filtering design relies solely on tracklet features, a promising direction for future work is to dynamically filter noisy frames at different stages of training.